OpenStudy (anonymous):

. Line p passes through the point (5, 1) and has a slope of -4/5 . What is the equation of line p in standard form? 4x + 5y = 25 4x + 5y = 29 5x + 4y = 25 5x + 4y = 29

OpenStudy (anonymous):

use point-slop form y-y1 =m(x-x1) since you are given a point and the slope just plug it in to the equation

OpenStudy (anonymous):

so y-1=-4/5 (x-5) then simplify